OpenStudy (anonymous):

Solve for x: |x − 2| + 10 = 12

OpenStudy (anonymous):

This is a practical way to solve equations with absolute expression |x - 2| = 2 Solve the 2 equations: -(x - 2) = 2 (x - 2) = 2 -x + 2 = 2 --> x = 0 x = 4

OpenStudy (anonymous):

I think it's x=4,0 But I'm not 100% sure.